john earns $150,000 through his job as an accountant in 2019. What is his FICA tax as an employee?
Answer #1

Answer :- $10,414.80

- FICA TAX is a combination of Social Security tax and Medicare tax
- Income upto $132,900 will be subject to social security tax
Social Security tax rate = 6.2%
Medicare tax                  = 1.45%
Social Security Tax $132,900 * 6.2% = $8,239.80
Medicare    Tax         $150,000 * 1.45% = $2,175
Total FICA ( 8,239.80+2,175) = 10,414.80